OpenStudy (anonymous):

Solve the equation. –6p + 7 = 3(2p – 3) – 4(–10 + 4p)

OpenStudy (anonymous):

\[-67=6p-9+40-16p\] \[-67-40+9=6p-16p\] \[-98=-10p\] \[\frac{-98}{-10}=p\] \[9.8=p\]

OpenStudy (anonymous):

i think so !!! not sure

OpenStudy (anonymous):

only answers are 6, 5, 7, 17